# Practical Insights into Knowledge Distillation for Pre-Trained Models

### v1

@@ IV-A Different transfer set options
Figure 3: Optimization of hyperparameters $T$ and $\alpha$ in KD across various data partitioning strategies. The first column illustrates optimal $T$ values from grid search, related to performance differences between student and teacher models. The second column shows the best $\alpha$ values. The third column presents cumulative accuracy gains in percentage points from different KD methods. The final column details the accuracy gains for each participant pair using various KD techniques.
(a) Label Skew Random Chunks Distribution, CINIC10
Figure 4: Optimization of hyperparameters $T$ and $\alpha$ in KD for CINIC-10 dataset.
IV-C Optimizing hyperparameters: Tuned KD effectiveness
@@
Fig. 3 illustrates the main outcomes of our hyperparameter optimization endeavors. It highlights the impact of varying $T$ and $\alpha$ values on KD performance across different data partitioning strategies and identifies the optimal settings for these parameters.
Figure 5: Comparative analysis of the best teachers across diverse KD scenarios. The bar charts in the first row depict the frequency of each participant being selected as the best teacher within four different KD approaches, across five data distribution strategies. The heatmaps in the second row illustrate the corresponding participant accuracies, providing insight into the relationship between a participant’s accuracy and their likelihood of being an optimal teacher. The x-axis of the first row represents the participant’ IDs, which align directly above their respective accuracy heatmap.
@@
As demonstrated in the last column of Fig. 3, Tuned KD consistently outperforms Vanilla KD, DML, and DP-KD in all data distribution scenarios. It is particularly noteworthy that Tuned KD achieves positive knowledge transfers even with weak teachers, markedly differing from the outcomes seen with Vanilla KD and DP-KD. Additionally, in situations involving strong teachers, Tuned KD enhances accuracy gains, surpassing all other KD approaches. Furthermore, the performance analysis summarized in the third column of Fig. 3 depicts the cumulative accuracy gain achieved by each KD method across different data distributions, reinforcing the standout performance of Tuned KD. This finding underscores the importance of optimally tuning the $T$ and $\alpha$ parameters. It also opens up new avenues for future research, particularly in developing methods for efficient and automated tuning of these parameters tailored to the specific setting and learning objectives.
@@
The last column of Fig. 3 illustrates this phenomenon clearly. In scenarios where there is a significant performance disparity between the teacher and the student, Tuned KD demonstrates a marked improvement over Vanilla KD. Conversely, in cases where the teacher and student exhibit similar performance levels, the benefits of tuning $T$ and $\alpha$ diminish, resulting in Tuned KD achieving performance levels comparable to those of Vanilla KD.
@@
Fig. 3, particularly columns 1 and 2, illustrates the optimal $T$ and $\alpha$ values determined through our grid search across diverse data distributions. While identifying a consistent pattern for the best $T$ and $\alpha$ values in the Uniform Data Distribution is challenging, clear trends emerge in the Quantity Skew Distribution, the Label Skew Random Chunks Distribution, and the Label Skew Dirichlet Distribution. In these scenarios, a smaller $T$ value tends to be more effective with a weaker teacher, while a larger $T$ value is advantageous for a stronger teacher. Correspondingly, the ideal $\alpha$ value, indicating the weight assigned to the teacher’s output, should align with the teacher’s relative strength compared to the student. Thus, a significantly stronger teacher warrants a higher $\alpha$ value (above 0.5), while a weaker teacher should receive a lower weight. Moreover, the Specialized Non-IID Distribution generally favors lower values of $T$ and $\alpha$, leading to more favorable outcomes across all teacher-student pairings.
Fig. 4 includes experiments using the CINIC10 dataset [28], albeit with a singular data distribution approach, serves to validate and reinforce our findings from the CIFAR10 dataset. This consistency indicates result robustness and suggests the potential generalizability of our conclusions across different datasets, opening avenues for further exploration in diverse data environments.
@@ IV-D Best teachers across KD methods and data distributions
Fig. 5 juxtaposes the frequency of participants emerging as the best teachers against their respective performance within varied data distribution strategies, utilizing the four KD approaches. The best teacher is defined as one that provides the greatest accuracy gain for a given student participant. This analytical approach aims to discern the potential correlation between a participant’s accuracy and its propensity to be an effective teacher.
@@
The findings indicate that participant accuracy, though pivotal, is not the only factor influencing the selection of the best teacher across diverse KD techniques and data distributions. This highlights the complexity of the KD teacher selection process, which appears to consider a broader spectrum of factors.
Under the uniform data distribution, the emergence of the best teachers appears to be evenly spread across participants for most KD approaches. In contrast, in scenarios with significant data skewness, particularly in quantity and label skew distributions, KD methods such as Tuned KD, Vanilla KD, and DP-KD tend to prefer teachers with high accuracy, but also possibly incorporate other factors in their selection.
Interestingly, the DML approach’s teacher selection exhibits a weaker correlation with accuracy, implying that it may value additional attributes such as the diversity or specificity of participant knowledge.
This intricate selection dynamic indicates an opportunity to enhance KD algorithms by integrating a more complex array of criteria beyond accuracy. Subsequent studies should aim to decode these implicit factors, potentially leading to more refined and efficacious KD paradigms.
@@
IV-E Revisiting KD methods: A comparative perspective through data distributions and learning dynamics
In this subsection, we re-examine the performance of various KD methods, including Tuned KD, under a different light. This revisit emphasizes a comparative analysis grounded in different data distribution strategies, consolidating our earlier observations and unveiling new insights into the adaptability and robustness of these methods.
Fig. 6 juxtaposes the efficacy of KD methods against varied data distributions. This comparative visualization allows to discern patterns and outliers in performance, highlighting the consistent strength of Tuned KD across multiple scenarios. By contrasting these methods side-by-side, we gain a clearer understanding of how each method responds to diverse data conditions.
Fig. 7 investigates the learning-forgetting (cf. § III) trade-off in KD methods. Tuned KD and DML show less forgetting with weaker teachers, outperforming Vanilla KD and DP-KD. Uniform data distributions reveal similar performances across methods, with a minor uptick in forgetting for DP-KD. Tuned KD emerges as the most effective, maximizing learning in all scenarios and reducing forgetting in Quantity Skew and Specialized Non-IID distributions. In contrast, it shows increased forgetting within Label Skew distributions. DML shows minimal forgetting but is constrained in learning gains, suggesting a conservative knowledge consolidation preference. This analysis underscores the need to tailor KD approach selection to the specific strengths and weaknesses of each method, aligned with data characteristics and learning goals.
Finally, the decision tree in Fig. 8 serves as a strategic guide for selecting suitable KD methods based on specific use cases and data environments, offering a practical approach for applying KD methods in various contexts.
V KD from Multiple Pre-trained Models
This section explores the application of KD in the context of multiple pre-trained models. Our investigation primarily revolves around Vanilla KD, examining its effectiveness in consolidating knowledge from various models, each with distinct training experiences and data exposures. This study aligns with the broader goal of harnessing the collective intelligence of diverse models in decentralized and federated learning systems. Additionally, the section briefly touches upon the application of KD in enhancing FL, assessing its role in pre-consolidating models before initiating the FL process. This aspect of the study emphasizes the potential of KD in streamlining and advancing the FL workflow.
TABLE II: Combining multiple models across various data distributions and transfer set options
Data Distribution
Public Unlabeled Transfer Set Student Data as Transfer Set
@@ V KD from Multiple Pre-trained Models
Uniform
71.64 72.36 72.35 73.43 69.55 70.59 71.34 72.24 72.68 73.97
Quantity Skew
68.18 64.89 77.90 73.86 67.31 64.66 29.77 30.01 78.34 76.33
Specialized Non-IID
51.57 54.72 53.07 55.11 52.55 52.98 43.70 50.72 52.39 53.60
Label Skew Random Chunks
65.44 66.73 73.33 71.68 66.35 66.22 19.22 20.67 75.71 75.41
Label Skew Dirichlet
61.67 54.03 67.63 57.82 63.77 52.41 28.61 27.35 66.92 63.00
V-A Combining multiple models
We delve into methodologies for the effective amalgamation of knowledge across multiple pre-trained models via KD. Our focus includes critical aspects such as the selection of an appropriate starting student model and the assessment of various teacher weighting strategies across different transfer set scenarios.
Where to start?  Selecting the optimal starting student model is key in KD. We explore three scenarios: using the lowest-performing, highest-performing, or an untrained student model.
Transfer set.  Our approach incorporates two transfer set options: a public unlabeled dataset with 5,000 samples, and the student’s dataset. The models undergo 30 epochs of training over these transfer sets.
Teachers’ weights.  We investigate equal weighting for all teachers and adaptive weighting, which is based on each teacher’s class-wise contribution.
Here, $J$ is the total number of teachers, and the dimension of each $\alpha_{j}$ vector matches the number of classes.
The $\mathcal{L}_{KL}$ in Eq. 2 can be updated considering the combined influence of multiple teachers, as:
Table II presents a summary of our experimental outcomes, offering insights into the efficacy of different KD strategies under varying conditions. The ‘Weighted’ and the ‘Unweighted’ columns in the table represent the adaptive weighting and the equal weighting strategies, respectively.
@@ V-A Combining multiple models
The data suggest a consistent trend: initiating the distillation process with the best-performing student model generally leads to better results. This could be attributed to the more advanced starting point it provides, requiring less knowledge to assimilate.
Finding (11): The adaptive weighting approach shows varied effectiveness based on data distribution, with marked gains in Label Skew Dirichlet and Quantity Skew scenarios.
The adaptive weighting approach delineated in Eq. 6 and Eq. 7 manifests considerable enhancements in specific data distribution contexts, notably the Label Skew Dirichlet and Quantity Skew scenarios. Conversely, in the Specialized Non-IID and the Uniform data distributions, the equal weighting strategy across all teachers appears to be more beneficial. This variability underscores the need to tailor weighting strategies based on specific data contexts in KD applications.
@@
However, the comparison between using the best student’s data and the worst student’s data as the transfer set shows a marked difference. Opting for the best student’s data typically leads to significantly better results. This can be attributed to the broader and possibly richer data representation in the best student’s dataset, in contrast to the more limited scope of the worst student’s data. This finding highlights the importance of transfer set size and quality, corroborating the initial observations we made in § IV regarding the impact of transfer set on KD performance.
In light of these observations, we will proceed with further experiments using the student data as the transfer set and employing the adaptive weighting mechanism for knowledge distillation.
V-B Accelerated FL with KD pre-consolidation
This subsection explores the application of KD to enhance FL. We assess the effectiveness of using KD for model pre-consolidation before initiating FL, particularly in cross-silo settings, a domain directly relevant to our research.
We compare traditional FL using FedAvg with the proposed approach where models first undergo KD pre-consolidation. This method utilizes the pre-consolidated model as the initial model for FL, diverging from the standard practice of starting with randomly initialized models. Our analysis spans various data distribution scenarios to gauge the impact of KD pre-consolidation on reducing the number of communication rounds necessary to achieve target accuracy in FL.
Fig. 9 presents the results of this comparative analysis.
@@ V-B Accelerated FL with KD pre-consolidation
KD consolidation, when used alone, might not always produce satisfactory results in complex environments like Specialized Non-IID and Label Skew Dirichlet distributions. However, incorporating it as a preliminary step in FL significantly speeds up the learning process. This demonstrates the potential of KD pre-consolidation as a powerful tool for improving the efficiency of FL systems, particularly in accelerating the convergence towards target accuracy levels.
VI Conclusion
This paper presents an in-depth exploration of KD in the context of pre-trained models, offering vital insights and empirical evidence into its effectiveness. We have systematically compared various KD methodologies, such as Vanilla KD, Tuned KD, DML, and DP-KD, across multiple data distribution strategies, highlighting their unique strengths and applications. Our findings demonstrate that Tuned KD, particularly with optimized parameters, consistently outperforms other methods in various scenarios.
Furthermore, we have unveiled the potential of KD in accelerating FL. By pre-consolidating models using KD, our experiments indicate a significant reduction in the number of communication rounds required for FL, thus enhancing its efficiency.
Overall, our research enriches the understanding of KD within distributed learning systems. It underscores KD’s strategic value in enhancing both model performance and the efficiency of training processes, setting the stage for further advancements in both centralized and decentralized distributed training environments.

### v2

@@ IV-C Optimizing Hyperparameters: Tuned KD
Refer to caption Refer to caption              Refer to caption Refer to caption
@@
Refer to caption Refer to caption          Refer to caption Refer to caption
@@
Refer to caption Refer to caption              Refer to caption Refer to caption
@@
Refer to caption Refer to caption               Refer to caption Refer to caption
@@
Figure 4: Optimization of hyperparameters $T$ and $\alpha$ in KD across various data partitioning strategies for the CIFAR-10 dataset. The first column illustrates the best $T$ values derived from grid search, related to performance differences between student and teacher models. The second column shows the best $\alpha$ values. The third column presents cumulative accuracy gains in percentage points from different KD methods. The final column details the accuracy gains for each participant pair using various KD techniques.
Refer to caption Refer to caption        Refer to caption Refer to caption
Figure 5: Optimization of hyperparameters $T$ and $\alpha$ in KD for CINIC-10 dataset using Label Skew Random Chunks Distribution.
@@
Fig. 4 illustrates the main outcomes of our hyperparameter optimization endeavors. It highlights the impact of varying $T$ and $\alpha$ values on KD performance across different data partitioning strategies and identifies appropriate settings for these parameters.
@@
As demonstrated in the last column of Fig. 4, Tuned KD consistently outperforms Vanilla KD, DML, and DP-KD in all data distribution scenarios. It is particularly noteworthy that Tuned KD achieves positive knowledge transfers even with weak teachers, markedly differing from the outcomes seen with Vanilla KD and DP-KD. Tuned KD enhances accuracy gains in situations involving strong teachers, surpassing all other KD approaches. Furthermore, the performance analysis summarized in the third column of Fig. 4 depicts the cumulative accuracy gain achieved by each KD method across different data distributions, reinforcing the standout performance of Tuned KD. This finding underscores the importance of optimally tuning the $T$ and $\alpha$ parameters. It also opens up new avenues for future research, particularly in developing methods for efficient and automated tuning of these parameters tailored to the specific setting and learning objectives.
Figure 6: Comparative analysis of the best teachers across diverse KD scenarios. The bar charts in the first row depict the frequency of each participant being selected as the best teacher within four different KD approaches across five data distribution strategies. The heatmaps in the second row illustrate the corresponding participant accuracies, providing insight into the relationship between a participant’s accuracy and their likelihood of being an optimal teacher. The x-axis of the first row represents the participant’ IDs, which align directly above their respective accuracy heatmap.
@@
The last column of Fig. 4 illustrates this phenomenon clearly. In scenarios with a significant performance disparity between the teacher and the student, Tuned KD demonstrates a notable improvement over Vanilla KD. Conversely, in cases where the teacher and student exhibit similar performance levels, the benefits of tuning $T$ and $\alpha$ diminish, resulting in Tuned KD achieving performance comparable to Vanilla KD’s.
@@
Fig. 4, particularly columns 1 and 2, illustrates the optimal $T$ and $\alpha$ values determined through our grid search across diverse data distributions. While identifying a consistent pattern for the best $T$ and $\alpha$ values in the Uniform Data Distribution is challenging, clear trends emerge in the Quantity Skew Distribution, the Label Skew Random Chunks Distribution, and the Label Skew Dirichlet Distribution. In these scenarios, a smaller $T$ value tends to be more effective with a weaker teacher, while a larger $T$ value is advantageous for a stronger teacher. Correspondingly, the ideal $\alpha$ value, indicating the weight assigned to the teacher’s output, should align with the teacher’s relative strength compared to the student. Thus, a significantly stronger teacher warrants a higher $\alpha$ value (above 0.5), while a weaker teacher should receive a lower weight. Moreover, the Specialized Non-IID Distribution favors lower values of $T$ and $\alpha$, leading to more favorable outcomes across all teacher-student pairings.
Fig. 5 shows the results from the experiments using an alternative practical CINIC-10 dataset [31] to validate and reinforce our findings from the CIFAR-10 dataset. In these experiments, we opt for the challenging ‘Label Skew Random Chunks’ data distribution approach. The results indicate the consistency of our observations drawn from CIFAR-10 results and suggest the potential generalizability of our conclusions about the efficacy of tuning across different benchmarks, especially in diverse data environments.
@@ IV-D Best teachers across KD methods and data distributions
Fig. 6 juxtaposes the frequency of participants emerging as the best teachers against their respective performance within varied data distribution strategies, utilizing the four KD approaches. The best teacher is defined as one who provides the greatest accuracy gain for a given student participant. This analytical approach aims to discern the potential correlation between a participant’s accuracy and their propensity to be an effective teacher.
@@
The findings indicate that participant accuracy, though pivotal, is not the only factor influencing the selection of the best teacher across diverse KD techniques and data distributions. This highlights the complexity of the KD teacher selection process, which considers a broader spectrum of factors.
Under the uniform data distribution, the emergence of the best teachers appears to be evenly spread across participants for most KD approaches. In contrast, in scenarios with significant data skewness, particularly in quantity and label skew distributions, KD methods such as Tuned KD, Vanilla KD, and DP-KD tend to prefer highly accurate teachers, but also possibly incorporate other factors in their selection.
Interestingly, the DML approach’s teacher selection exhibits a weaker correlation with accuracy, implying that it may value additional attributes such as the diversity or specificity of participant knowledge. This intricate selection dynamic indicates an opportunity to enhance KD algorithms by integrating a more complex array of criteria beyond accuracy. Subsequent studies should aim to decode these implicit factors, potentially leading to more refined and efficacious KD paradigms.
IV-E Model Learning and Forgetting Dynamics
We examine the relation between KD methods and the model’s learning and forgetting under different data distribution strategies, consolidating our earlier observations and unveiling new insights into the adaptability and robustness of these methods.
Fig. 7 investigates the learning-forgetting (cf. § III) trade-off in KD methods. Tuned KD and DML show less forgetting with weaker teachers, outperforming Vanilla KD and DP-KD. Uniform data distributions reveal similar performances across methods, with a minor uptick in forgetting for DP-KD. Tuned KD emerges as the most effective, maximizing learning in all scenarios and reducing forgetting in Quantity Skew and Specialized Non-IID distributions. In contrast, it shows increased forgetting within Label Skew distributions. DML shows minimal forgetting but is constrained in learning gains, suggesting a conservative knowledge consolidation preference. This analysis underscores the need to tailor the KD approach selection to each method’s specific strengths and weaknesses and align it with data characteristics and learning goals.
@@ IV-E Model Learning and Forgetting Dynamics
IV-F Influence of Data Distribution and Design Choices for KD
We re-examine the performance of various KD methods, including Tuned KD, under a different light and unveil new insights into the adaptability and robustness of these methods. Fig. 8 juxtaposes the efficacy of KD methods against varied data distributions. This comparative visualization allows us to discern patterns and outliers in performance, highlighting the consistent strength of Tuned KD across multiple scenarios. By contrasting these methods side-by-side, we better understand how each method responds to diverse data conditions.
{mdframed}
[backgroundcolor=white, linecolor=gray, innertopmargin=7pt, innerbottommargin=7pt, innerleftmargin=7pt, innerrightmargin=7pt, roundcorner=2pt] Refer to caption
@@ IV-F Influence of Data Distribution and Design Choices for KD
Finally, based on the previous findings, we create the decision tree shown in Fig. 9 as a strategic guide for selecting suitable KD methods based on specific use cases and data environments, offering a practical approach for applying KD methods in various contexts.
TABLE II: Results of combining multiple models across various data distributions and transfer set options. The table presents accuracy in percentage points post-distillation.
Data Distribution Public Unlabeled Transfer Set Student Data as Transfer Set
@@
Uniform 71.64 72.36 72.35 73.43 69.55 70.59 71.34 72.24 72.68 73.97
Quantity Skew 68.18 64.89 77.90 73.86 67.31 64.66 29.77 30.01 78.34 76.33
Specialized Non-IID 51.57 54.72 53.07 55.11 52.55 52.98 43.70 50.72 52.39 53.60
Label Skew Random Chunks 65.44 66.73 73.33 71.68 66.35 66.22 19.22 20.67 75.71 75.41
Label Skew Dirichlet 61.67 54.03 67.63 57.82 63.77 52.41 28.61 27.35 66.92 63.00
IV-G KD from Multiple Pre-trained Models
This subsection evaluates the application of KD in the context of multiple pre-trained models. Our investigation primarily revolves around Vanilla KD, examining its effectiveness in consolidating knowledge from various models, each with distinct training experiences and data exposures.
This analysis aligns with the broader goal of harnessing the collective intelligence of diverse models in centralized collaborative learning and FL systems.
Additionally, we examine how KD can be strategically employed to pre-consolidate knowledge across models before they participate in the FL process. By doing so, we aim to streamline the FL workflow, potentially reducing the number of communication rounds needed and accelerating the convergence toward optimal model performance.
IV-G1 Combining multiple models
We here delve into methodologies for the effective amalgamation of knowledge across multiple pre-trained models via KD. Our focus includes critical aspects such as the selection of an appropriate starting student model and the assessment of various teacher weighting strategies across different transfer set scenarios.
Where to start?  Selecting the starting student model is crucial in knowledge distillation. We explore three scenarios to determine the most effective approach for initiating the student model for the amalgamation process: 1) the worst-performing model among the group; 2) the best-performing model; or 3) an untrained student model, which begins from a completely randomized, untrained state.
Transfer set.  Our approach incorporates two transfer set options: 1) a public unlabeled dataset with 5,000 samples; and 2) the student’s dataset. The models undergo 30 epochs of training over these transfer sets.
Figure 10: FL Process with Pre-Consolidated Model Initialization. The server initializes the global model using the pre-consolidated model (Step 0), which incorporates knowledge from multiple teacher models. Step 0 occurs only once at the beginning. The subsequent steps (1-4) follow the standard FL process: the global model is sent to clients (Step 1), updated locally by each client (Step 2), sent back to the server (Step 3), and aggregated to form a new global model (Step 4). This process is repeated for a predefined number of rounds or until a certain performance criterion is met.
Table II summarises our experimental outcomes, offering insights into the efficacy of different KD strategies under varying conditions. The ‘Weighted’ and the ‘Unweighted’ columns in Table II represent the adaptive weighting and the equal weighting strategies, respectively (cf. Eq. 7).
@@ IV-G1 Combining multiple models
The data show a consistent trend: initiating the distillation process with the best-performing student model leads to better results. This can be due to its own developed knowledge.
Finding (11): The adaptive weighting approach shows varied efficacy based on data distribution, with marked gains in Label Skew Dirichlet and Quantity Skew scenarios.
The adaptive weighting approach delineated in Eq. 6 and Eq. 7 manifests considerable enhancements in specific data distribution contexts, notably the Label Skew Dirichlet and Quantity Skew scenarios. Conversely, the equal weighting strategy across all teachers appears more beneficial in the Specialized Non-IID and the Uniform data distributions. This variability underscores the need to tailor weighting strategies based on specific data contexts in KD applications.
@@
However, comparing using the best and worst student data as the transfer set shows a marked difference. Opting for the best student data typically leads to significantly better results. This can be attributed to the broader and possibly richer data representation in the best student’s dataset, in contrast to the more limited scope of the worst student’s data. This finding highlights the importance of transfer set size and quality, corroborating the initial observations we made in § IV regarding the impact of transfer set on KD performance.
Given these observations, we proceed with further experiments using the student data as the transfer set and employ the adaptive weighting mechanism for knowledge distillation.
IV-G2 Accelerated FL with KD pre-consolidation
We introduce a novel method of using KD for model pre-consolidation to enhance centralized collaborative training, particularly in cross-silo FL settings. Our approach deviates from traditional FL practices that typically begin with randomly initialized models. Instead, it involves a preliminary phase where multiple models from different silos are first consolidated using KD, as depicted in Fig. 2. This pre-consolidation process entails distilling knowledge from various pre-trained models into a single, unified model. This pre-consolidated model serves as the starting point for federated learning (Fig. 10), aiming to reduce the heterogeneity typically encountered at the onset of traditional FL.
We compare traditional FL using FedAvg against our proposed approach, which utilizes the pre-consolidated model as the initial model for FL. Our analysis spans various data distribution scenarios to gauge the impact of KD pre-consolidation on reducing the number of communication rounds necessary to achieve target accuracy in FL. Fig. 11 presents the results of this comparative analysis.
@@ IV-G2 Accelerated FL with KD pre-consolidation
When used alone, KD consolidation might not always produce satisfactory results in complex environments like Specialized Non-IID and Label Skew Dirichlet distributions. However, incorporating it as a preliminary step in FL significantly speeds up the learning process. This demonstrates the potential of KD pre-consolidation as a powerful tool for improving the efficiency of FL systems, particularly in accelerating the convergence toward target accuracy levels.
V Conclusion
This paper presents a detailed examination of knowledge distillation within collaborative learning contexts, particularly focusing on scenarios that employ pre-trained models. Our in-depth exploration provided crucial insights into optimizing KD techniques for enhancing collaborative learning effectiveness.
We have systematically compared various KD methodologies, such as Vanilla KD, Tuned KD, DML, and DP-KD, across multiple data distribution strategies, highlighting their unique strengths and applications. Our findings demonstrate that Tuned KD consistently outperforms other methods in various scenarios, particularly with optimized parameters.
Our research reveals the potential of KD to accelerate FL in the context of centralized collaborative learning systems. By pre-consolidating models using KD, our experiments indicate a substantial reduction in the number of communication rounds required for FL, thus enhancing its efficiency. Overall, our research enriches the understanding of KD within collaborative learning systems.